# FedVSR: Towards Model-Agnostic Federated Learning in Video Super-Resolution

### v2

@@ 4.1.1 DWT for High-Frequency Preservation
As our experiments demonstrate (see Section 5.3), the optimal weight values are unbounded, meaning they are not constrained by the relationship $\lambda_{\text{VSR}}=1-\lambda_{\text{HiFr}}$ and can be independently adjusted for better performance.
4.2 Loss-Aware Aggregation for Federated VSR
To further enhance the model’s performance, we employ a greedy aggregation approach with a decay rate, inspired by MKL-SGD [46]. MKL-SGD improves robustness by selecting samples with the lowest loss during training, thereby mitigating the impact of outliers. Similarly, in our method, clients accumulate their loss values throughout local training and compute an average loss based on the number of training steps. When sending their updated model weights to the server, they also transmit this average loss. At the server, we perform inverse weighted averaging, where each client’s weight is computed as $I_{i}=\frac{1}{\mathcal{L}_{i}}$ and adjusted using an adaptive step: $I^{\prime}_{i}=I_{i}^{\alpha}$ where $\alpha$ controls the strength of loss-based weighting. To regulate this process, we introduce a decay rate that gradually reduces the effect of loss-based weighting over time, following:$\alpha=\alpha^{1-\frac{t}{T}}$ where $t$ is the current round and $T$ is the total number of rounds. This ensures that early training benefits from loss-based prioritization, while later rounds transition smoothly toward standard averaging, balancing robustness and generalization. The overall algorithm is summarized in Algorithm 1.
Algorithm 1 Our Method
1:Number of communication rounds $T$, Number of clients $N$, Learning rate $\eta$, Adaptive step $\alpha$, Decay rate $\gamma$
2:Global model weights $W_{T}$
3:Server Initialization Initialize global model $W_{0}$
@@ 4.2 Loss-Aware Aggregation for Federated VSR
7:        Client updates local model:
8:         Retrieve latest global model $W_{t}$
9:         Perform local training for $E$ epochs with learning rate $\eta$
10:        Compute Local Loss:
11:         Compute model-specific loss $\mathcal{L}_{\text{VSR}}$
12:         Compute $\mathcal{L}_{\text{DWT}}$ loss $\mathcal{L}_{\text{DWT}}$
13:         Compute total loss
$\mathcal{L}=\lambda_{\text{VSR}}\mathcal{L}_{\text{VSR}}+\lambda_{\text{HiFr}}\mathcal{L}_{\text{HiFr}}$
14:        Client updates local model:
$\mathbf{w}_{i}^{t+1}=\mathbf{w}^{t}-\eta\nabla F_{i}(\mathbf{w}^{t})$
15:         Return updated weights $W_{i}$ and loss value $\mathcal{L}_{i}$ to server
16:    end for
17:    Server Aggregation: (Performed at the Server)
18:     Compute inverse loss weighting $I_{i}=\frac{1}{\mathcal{L}_{i}}$
19:     Apply adaptive step $I^{\prime}_{i}=I_{i}^{\alpha}$
20:     Normalize inverse loss weights $w_{i}=\frac{I^{\prime}_{i}}{\sum_{j\in S_{t}}I^{\prime}_{j}}$
21:     Compute weighted model update $W_{i}^{\text{scaled}}=w_{i}\cdot W_{i}$
22:     Aggregate global model $W_{t+1}=\sum_{i\in S_{t}}W_{i}^{\text{scaled}}$
23:    Apply Decay Rate to Adaptive Step $\alpha=\alpha^{1-\frac{t}{T}}$
24:end for
25:return Global model $W_{T}$
@@ 5 Experiments
Model Algorithm Less Heterogeneous More Heterogeneous

### v3

@@ 4.1 High-Frequency Loss for Federated VSR
4.2 Adaptive Loss-Aware Aggregation
To further enhance the model’s performance, we employ a hybrid aggregation strategy that dynamically balances uniform and loss-based weighting.
In general VSR research, the standard task loss (e.g., L1 or L2) is already widely accepted as a reliable pixel-wise measure of reconstruction quality [28]. This provides a crucial distinction from federated classification tasks. The total local loss ($\mathcal{L}_{\text{total}}$) used in FedVSR maintains this reliability, being calculated as the combination of the standard VSR loss and our novel $\mathcal{L}_{\text{HiFr}}$ high-frequency loss. Our $\mathcal{L}_{\text{HiFr}}$ term acts as an enhancement, sensitizing $\mathcal{L}_{\text{total}}$ to the high-frequency details often lost through averaging, thus making the total loss an even better measure of performance. Unlike high-level tasks such as classification where a high loss might signal valuable Non-IID data [53] , a high value for $\mathcal{L}_{\text{total}}$ in VSR directly signifies a failure to reconstruct high-fidelity details and temporal coherence, the fundamental objective. Therefore, clients with significantly higher $\mathcal{L}_{\text{total}}$ pixel-wise losses genuinely contribute lower-quality model updates, making prioritization justifiable.
This aggregation strategy, inspired by MKL-SGD [40], improves robustness by prioritizing updates that achieve superior local performance. Our method first computes the average local loss ($\mathcal{L}_{\text{total}}$) for each client after training, which is then transmitted to the server as a single scalar value alongside the model weights.
4.2.1 Deriving Loss-Based Prioritization
Rather than using inverse loss weights directly, we normalize them and apply an adaptive step parameter to create the loss-based distribution ($l_{i}$), which naturally emphasizes clients with lower losses.
4.2.2 Hybrid Weighting for Robustness
A key design feature is that we do not rely on $l_{i}$ alone. The aggregation is governed by a convex combination of $l_{i}$ and the uniform distribution ($u_{i}$). This $\mathbf{u_{i}}$ component ensures that all clients contribute a baseline minimum update, thereby preventing the complete marginalization of updates derived from challenging, yet strategically important, diverse video data.
4.2.3 Thresholded Adaptation via Hellinger Distance
To avoid overreacting to minor variations in client performance, we quantify the divergence between the uniform ($u_{i}$) and loss-based ($l_{i}$) distributions using the Hellinger distance [16]. This distance serves as a mixing coefficient ($m$). By applying a soft threshold ($\tau$) to this distance, the loss-based prioritization is only significantly activated when there is a large, meaningful divergence in client update quality. If losses are similar, the aggregation gradually defaults to uniform averaging.
Finally, the client’s weight ($w_{i}$) in the global update is the convex combination of these two distributions, which is then applied to the model parameters during aggregation. This approach allows the aggregation process to adaptively leverage loss-based prioritization when needed, while safeguarding against the unfair penalization of diverse data, thereby improving both robustness and generalization. The overall algorithm is summarized in Algorithm 1.
Algorithm 1 Our Method (FedVSR)
1:Number of communication rounds $T$, Number of clients NN, Learning rate $\eta$, Adaptive step $\alpha$, Decay rate $\gamma$
2:Global model weights $W^{T}$
3:Server Initialization: Initialize global model $W^{0}$
@@ 4.2.3 Thresholded Adaptation via Hellinger Distance
7:    Client downloads latest global model $W^{t}$
8:    Perform local training for $E$ epochs with learning rate $\eta$
9:    Compute local losses:
10:       $\mathcal{L}_{\text{VSR}}$ (task-specific loss)
11:       $\mathcal{L}_{\text{HiFr}}$ (high-frequency loss)
12:    Compute total loss:
$\mathcal{L}_{i}=\lambda_{\text{VSR}}\mathcal{L}_{\text{VSR}}+\lambda_{\text{HiFr}}\mathcal{L}_{\text{HiFr}}$
13:    Update local model:
$W_{i}^{t+1}=W^{t}-\eta\nabla F_{i}(W^{t})$
14:    Return $W_{i}^{t+1}$ and loss value $\mathcal{L}_{i}$ to server
15:  end for
16:  Server Aggregation:
17:  Compute uniform weights $u_{i}=\frac{1}{|S_{t}|}$
18:  Compute inverse-loss weights $I_{i}=\frac{1}{\mathcal{L}_{i}}$, apply adaptive step $I^{\prime}_{i}=I_{i}^{\alpha}$
19:  Normalize loss-based weights $l_{i}=\frac{I^{\prime}_{i}}{\sum_{j\in S_{t}}I^{\prime}_{j}}$
20:  Compute Hellinger distance:
$H=\sqrt{\tfrac{1}{2}\sum_{i\in S_{t}}(\sqrt{u_{i}}-\sqrt{l_{i}})^{2}}$
21:  Compute mixing coefficient:
$m=\begin{cases}0&H<\tau\\ \tfrac{H-\tau}{1-\tau}&H\geq\tau\end{cases}$
22:  Final aggregation weights: $w_{i}=(1-m)u_{i}+ml_{i}$
23:  Update global model:
$W^{t+1}=\sum_{i\in S_{t}}w_{i}\cdot W_{i}^{t+1}$
24:  Apply Decay Rate to Adaptive Step $\alpha=\alpha^{1-\frac{t}{T}}$
25:end for
26:return Global model $W^{T}$
@@ 5 Experiments
Model Params (M) Algorithm Less Heterogeneous
REDS [35] Vid4 [26]
[…]
VRT [24] 35.6 FedAvg [30] 29.71 0.8522 0.1916 96.59 24.88 0.7680 0.2446 81.62
FedProx [21] 28.82 0.8281 0.2133 92.95 24.18 0.7341 0.2668 76.03
SCAFFOLD [18] 28.59 0.8181 0.2228 91.45 23.99 0.7182 0.2737 73.15
FedMedian [52] 27.89 0.7999 0.2379 87.46 23.59 0.6997 0.2963 70.62
FedVSR (Ours) 29.96 0.8595 0.1807 97.47 24.97 0.7745 0.2284 83.07
RVRT [23] 10.8 FedAvg [30] 31.30 0.8877 0.1536 97.97 25.96 0.8156 0.1974 86.05
FedProx [21] 30.09 0.8618 0.1899 95.65 25.20 0.7775 0.2330 82.01
SCAFFOLD [18] 29.63 0.7967 0.2101 94.71 24.79 0.7317 0.2407 81.01
FedMedian [52] 29.76 0.8444 0.1931 94.79 24.88 0.7602 0.2404 80.11
FedVSR (Ours) 31.64 0.8956 0.1377 98.94 26.10 0.8207 0.1802 87.37
IART [48] 13.4 FedAvg [30] 30.56 0.8729 0.1695 97.84 25.43 0.7949 0.2118 84.34
FedProx [21] 29.65 0.8500 0.1683 96.14 24.72 0.7599 0.2307 80.65
SCAFFOLD [18] 30.29 0.8678 0.1692 97.47 25.18 0.7857 0.2084 83.38
FedMedian [52] 29.17 0.8265 0.1960 95.55 24.40 0.7447 0.2302 78.50
FedVSR (Ours) 30.98 0.8845 0.1473 98.45 25.59 0.8077 0.1899 86.12
Table 1: Comparison of PSNR ($\uparrow$), SSIM ($\uparrow$), LPIPS ($\downarrow$), and VMAF ($\uparrow$) for different FL algorithms under less heterogeneous data distributions across VSR models and datasets.
Model Params (M) Algorithm More Heterogeneous
REDS [35] Vid4 [26]
[…]
VRT [24] 35.6 FedAvg [30] 28.85 0.8282 0.2200 91.81 24.60 0.7516 0.2557 77.52
FedProx [21] 28.64 0.8236 0.2222 88.04 24.10 0.7256 0.2713 72.16
SCAFFOLD [18] 28.17 0.8074 0.2318 87.83 24.04 0.7208 0.2739 71.85
FedMedian [52] 27.68 0.7761 0.2429 85.72 23.77 0.7001 0.2890 69.69
FedVSR (Ours) 29.12 0.8371 0.2010 94.37 24.78 0.7646 0.2378 80.10
RVRT [23] 10.8 FedAvg [30] 29.89 0.8578 0.1867 96.31 25.67 0.8040 0.2065 84.13
FedProx [21] 29.59 0.8526 0.1943 94.71 25.02 0.7730 0.2353 82.01
SCAFFOLD [18] 26.71 0.7142 0.2590 72.38 23.08 0.6441 0.3018 57.54
FedMedian [52] 28.85 0.8202 0.2119 87.84 24.51 0.7419 0.2499 74.73
FedVSR (Ours) 31.00 0.8844 0.1492 97.45 25.95 0.8197 0.1823 85.85
IART [48] 13.4 FedAvg [30] 29.52 0.8493 0.1909 95.99 25.04 0.7774 0.2272 81.71
FedProx [21] 29.30 0.8444 0.1958 94.16 24.42 0.7505 0.2293 77.42
SCAFFOLD [18] 29.26 0.8413 0.1992 94.16 24.85 0.7690 0.2278 79.40
FedMedian [52] 28.80 0.8280 0.2330 88.90 23.88 0.7350 0.2555 73.99
FedVSR (Ours) 29.85 0.8568 0.1787 96.84 25.31 0.7914 0.2118 83.18
Table 2: Comparison of PSNR ($\uparrow$), SSIM ($\uparrow$), LPIPS ($\downarrow$), and VMAF ($\uparrow$) for different FL algorithms under more heterogeneous data distributions across VSR models and datasets.
5.1 setup
5.1.1 Dataset:
We designed two distinct training environments to comprehensively evaluate the robustness and generalization of FedVSR under varying levels of data heterogeneity, utilizing two foundational video datasets: REDS [35] and Kinetics-400 [6]. The REDS dataset is high-quality and professionally recorded, designed specifically for video restoration tasks, offering high content diversity but remaining relatively controlled, with sequences at $720\times 1280$ resolution. In contrast, Kinetics-400,dataset primarily used for human action recognition, is a massive dataset sourced from unique amateur YouTube videos, whose content is highly realistic and exhibits significant camera shake, lighting variations, and low control, making it ideal for simulating highly diverse, uncontrolled, edge-device data.
For the Less Heterogeneous Scenario, we utilize the training sequences of REDS, splitting the data randomly ($\mathbf{IID}$-like) among clients to simulate a less diverse video environment. For the More Heterogeneous Scenario, we utilize a custom split of Kinetics-400 to simulate a strong $\mathbf{Non\text{-}IID}$ environment. Critically, because VSR does not use labels and developing an natural metric for video heterogeneity would be out of scope, we exploit the existing human action labels (40 classes selected) of Kinetics-400 solely to partition the video clips into Non-IID client silos, maximizing the distribution shift between clients. Only $720\times 1280$ videos are selected from Kinetics-400 for format consistency with REDS.
To assess the generalization capability of the models trained under both scenarios, we use two standard external test sets: the REDS Test set and the Vid4 dataset [26]. The models trained on both the REDS and Kinetics-400 splits are consistently evaluated against both benchmarks. The reason for selecting these specific datasets for testing is to strictly adhere to VSR community standards for better comparison of performance and measurement, since they are established benchmarks.
5.1.2 Implementation Details:
To maintain our objective of developing a model-agnostic framework, we did not alter the official parameters of the baseline VSR models. Instead, we used their recommended settings for training. We created 40 clients, with participation of 10% in both experimental setups. In the more heterogeneous setup, each client received videos from a single class; for example, one client had only ”archery” videos, while another had only ”making pizza” videos. In contrast, in the less heterogeneous setup, the six videos were randomly assigned to each client. We set the total number of global rounds to 100, with each client performing one local round (epoch) per global round. After completing their local training, these clients sent their updated weights to the server for aggregation. All experiments were conducted using bi-cubic (BI) degradation.
5.1.3 Baselines:
For our baseline, we use a combination of FedAvg [30], FedProx [21], FedMedian [52], and SCAFFOLD [18] from the FL side, along with VRT [24], RVRT [23], and IART [48] from VSR. We selected these FL methods as they are widely accepted and frequently used to benchmark core algorithmic performance in recent literature.
5.1.4 Evaluation Metrics:
The evaluation metrics used for the experiments are Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM) and Learned Perceptual Image Patch Similarity (LPIPS) and Video Multimethod Assessment Fusion (VMAF).
5.1.5 Reproducibility:
To ensure our experiments are reproducible, we fixed the random seed, which doesn’t affect client selection, model initialization, etc. We used four NVIDIA A6000 GPUs for training, with each model under a federated learning algorithm taking an average 46 hours to train. For the baselines alone, this consumed 1102 hours (2 experiments × 3 VSR models × 4 Fl algorithms × 46 hours per run).
5.2 Results and Comparison
5.2.1 Quantitative Results:
As shown in Tables  1 and  2, our proposed FedVSR algorithm consistently outperforms other federated learning (FL) algorithms across all tested models and datasets. On average, FedVSR surpasses existing FL methods by a margin of 0.82 dB in PSNR and 0.0337 in SSIM and 0.0334 in LPIPS and 3.62 in VMAF in less heterogeneous settings, while in more heterogeneous settings, the margins are 0.96 dB in PSNR and 0.0404 in SSIM and 0.0359 in LPIPS and 6.34 in VMAF, demonstrating its robustness under challenging conditions. Among the evaluated models, RVRT benefits the most from FedVSR, achieving the highest performance margin. In contrast, IART shows the lowest overall improvement but still maintains a competitive edge over the baseline FL methods.
The observed performance trends in our experiments align with existing findings on FedProx, SCAFFOLD and FedMedian further validating our results. As shown in Fig. 3, FedVSR consistently achieves superior PSNR values across different VSR models and datasets, outperforming FL baselines particularly in later training rounds. This aligns with prior research on FedProx, which highlights its failure to reach the stationary points of the global optimization objective, even under homogeneous data distributions. The study suggests that FedProx’s limiting points may remain far from the true minimizer of the global empirical risk [42], which could explain its suboptimal performance in our experiments. Although FedProx may still generalize well by aggregating local data effectively in certain conditions, its weaker convergence properties become apparent in more heterogeneous scenarios. We also observe that FedProx is highly sensitive to the $\mu$ value during training. The original paper suggests $\{0.001,0.01,0.1,1\}$ as candidate values and finds that $\mu=1$ performs better in most cases. However, we notice that for any $\mu$ greater than 0.001, the VSR model struggles significantly with learning. This suggests that the relationship is more complex and requires further investigation and analysis.
Figure 3: PSNR across different rounds for various test sets under different settings for VRT [24], RVRT [23], and IART [48].
Similarly, our findings regarding SCAFFOLD’s performance in highly heterogeneous settings support previous research showing that it is very sensitive to differences in client data [43]. We observed significant drops and stagnation in our experiments, particularly with the RVRT model, which matches reports that even small changes in local data batches can cause large swings in performance [43]. SCAFFOLD struggles to train effectively in such diverse scenarios, leading to slower or stalled convergence, which is reflected in our results 3 where its PSNR fluctuates in highly heterogeneous scenarios. The increased variance in SCAFFOLD’s performance under skewed data distributions further supports our findings that it fails to adapt smoothly across different client datasets, reinforcing its limitations in federated VSR tasks. On top of this, SCAFFOLD requires roughly twice the communication overhead of other algorithms due to its control variates, which further limits its practicality. See Section 5.3.6
Although FedMedian demonstrates better stability than SCAFFOLD in our experiments, it clearly suffers from a lower convergence rate, which is consistent with prior findings. Prior work [59] shows that the median-based update is less statistically efficient than mean or FedAvg when each client has a limited local dataset. Similarly, it is shown that coordinate-wise median becomes increasingly inefficient as the model grows in size or complexity, making it sub-optimal in high-dimensional settings. Additional empirical studies [32] note that FedMedian often converges more slowly because it lacks explicit regularization and tends to produce more conservative updates.
5.2.2 Qualitative Results:
As showen in Fig.  1 clearly demonstrate the limitations of FedAvg, FedProx, and SCAFFOLD in accurately reconstructing fine details. While these methods reduce some pixelation from the LQ input, they all fail to recover the textures and sharp edges present in the ground-truth. FedAvg produces overly smooth results, blurring important structural details such as the wooden lattice and frame edges. FedProx exhibits similar blurring, struggling to restore high-frequency textures. SCAFFOLD, not only fails to produce a sharp reconstruction but also introduces noticeable artifacts, distorting the texture of the wooden lattice and making the output appear unnatural. In contrast, FedVSR delivers significantly sharper and more realistic reconstructions, clearly outperforming the other approaches.
5.3 Extended Evaluation and Test Cases
5.3.1 Additional Baseline: FedSR vs FedVSR
As discussed in Section 2, existing FL methods for the SISR task were designed for a different setting, providing detailed SR model architectures and introducing extra overhead, which places them in a different category than FedVSR. To better understand how our model compares, we tested FedSR with VRT under both heterogeneity settings. Since the original paper did not provide a repository, we implemented the code as faithfully as possible. The results are shown in Table 3.
Algorithm Less Heterogeneous More Heterogeneous